OpenStudy (anonymous):

How would one solve this problem? Use the dot product to determine whether the vectors are parallel, orthogonal, or neither. \[v = ^\sqrt{2} j, w = 4i\] A. orthogonal B. parallel C. neither

ganeshie8 (ganeshie8):

\(\large v = 0i + \sqrt{2} j\) \(\large w = 4i + 0j\)

ganeshie8 (ganeshie8):

knw how to take the dot product, given two vectors ?

OpenStudy (anonymous):

Unfortunately, no... How do I do that though?

ganeshie8 (ganeshie8):

multiply the i componnets multiply the j components add them both

ganeshie8 (ganeshie8):

\(\large v = 0i + \sqrt{2} j\) \(\large w = 4i + 0 j\) \(\large v.w = 0*4 + \sqrt{2}*0 = 0+ 0 = 0\)

ganeshie8 (ganeshie8):

since the dot product is 0, the vectors will be perpendicular

ganeshie8 (ganeshie8):

another name for perpendicular is orthogonal

ganeshie8 (ganeshie8):

see if that makes more or less sense...

OpenStudy (anonymous):

May I ask what would make the vectors parallel?

ganeshie8 (ganeshie8):

the fancy way of saying it is : the vectors are parallel when the ratio of components are proportional

ganeshie8 (ganeshie8):

\(\large v = a_1i + b_1 j\) \(\large w = a_2i + b_2 j\) are parallel if \[\large \dfrac{a_1}{a_2} = \dfrac{b_1}{b_2}\]

OpenStudy (anonymous):

Thank you very much ^-^